OpenStudy (anonymous):

For f(x)=|x^2-ax| Find the maximum value of the function in the interval [a,0].

OpenStudy (amoodarya):

in this interval x(x-a)<0 so the function is f(x)=|x^2-ax|=ax-x^2 f'(x)=a-2x f'=0 ---> x=a/2 f(a/2)=a^2/2-a^2/4=a^2/4 f(0)=f(a)=0 so max of f(x) in this interval is a^2/4

OpenStudy (anonymous):

thanks. can you please explain a little. How did you get "in this interval x(x-a)<0"

OpenStudy (amoodarya):